# Partial identification and unmeasured confounding with multiple treatments and multiple outcomes

### v2

@@ 4 Partial identification with additional constraints
Theorem 2 highlights that with a factor confounding assumption, we can bound the causal effects for all estimands in (9) in settings with multiple treatments and multiple outcomes. However, in some contexts, these bounds are too wide to be practically useful without additional assumptions. Fortunately, there are additional benefits to joint inference on the partial identification region for multiple estimands. By way of example, suppose that we are interested in two estimands, $\text{PATE}_{a_{1},t_{1},t_{2}}$ and $\text{PATE}_{a_{2},t_{1},t_{2}}$, which correspond to disctinct outcomes, $a_{1}$ and $a_{2}$ but the same treatment contrast ($t_{2}$ vs $t_{1}$). Then, reasoning about the joint partial identification, as opposed to the marginal regions in (9), can be advantageous because additional assumptions about the bias for one estimand can provide information about the bias for the other.
In this bivariate example, the degree to which assumptions on one estimand inform the other is driven by the inner product of $a_{1}^{\top}\Gamma$ and $a_{2}^{\top}\Gamma$. If these two vectors are orthogonal, then knowledge of the bias for one estimand provides very little information on the bias for the other (Figure 1, left). However, if these vectors are nearly collinear, then knowledge of one is very informative on the value of the other, which can lead to much smaller joint partial identification regions (Figure 1, right). Note also that while we have focused on the confounding for the outcomes, given by $a_{1}^{\top}\Gamma$ and $a_{2}^{\top}\Gamma$, similar intuition holds for treatments. If the two estimands have different treatment contrasts, given by $(t_{1}^{(1)},t_{2}^{(1)})$ and $(t_{1}^{(2)},t_{2}^{(2)})$, then an equally important quantity is the inner product of the two vectors $(t_{1}^{(1)}-t_{2}^{(1)})^{\top}B$ and $(t_{1}^{(2)}-t_{2}^{(2)})^{\top}B$. If these two vectors have a large inner product, then knowledge of the bias for one estimand is very informative for the bias of the other.
@@
Figure 1: Comparison of bivariate partial identification regions for the bias when either $\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle=0$ or $\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle<0$.
@@ 4.1 Sensitivity Bounds on treatment effects
In many applications, prior knowledge on a reasonable range of effect sizes is known and can be incorporated into the partial identification regions. For instance, in the application to air pollution (Section 6), we know that it is biologically implausible for pollution to have strong protective effects on health outcomes. In other cases, we may have strong prior knowledge about the largest plausible harmful effect. In our notation, a direct assumption about the magnitude of the causal effect on a single estimand takes the form
@@
If we make these assumptions for a number of estimands, we gain more information about $R$, and hence the bias for all other estimands. Analytic characterizations of the partial identification region under these constraints is a non-trivial problem and as such, we find approximate solutions using a rejection sampling algorithm to explore the range of possible causal effects satisfying (10). Specifically, we draw orthogonal matrices $R^{(h)}$ for $h=1,\dots,H$ uniformly on the Stiefel manifold of $m\times m$ orthogonal matrices, for large values of $H$. For each orthogonal matrix, $R^{(h)}$, we check whether the constraints in (10) are satisfied. Each matrix $R^{(h)}$ that satisfies the aforementioned criteria determines the bias for each estimand of interest, and we can explore the set of these matrices to determine plausible biases, and therefore partial identification regions.
@@ 4.2 Sensitivity Bounds on Partial R-squared Values
As an alternative to specifying sensitivity bounds directly on the treatment effects, we can instead specify bounds on the partial variance explained by confounders. In a single-treatment single-outcome analysis, Cinelli & Hazlett (2020) propose specifying upper bounds on both the partial fraction of outcome variance explained by confounders ($R^{2}_{Y\sim U|T}$) and the fraction of treatment variance explained by confounders ($R^{2}_{T\sim U})$. Here, we show how their approach can be incorporated into our multi-treatment multi-outcome setting. Throughout this section, we let $d=(t_{1}-t_{2})$ correspond to the treatment contrast of interest.
For a given treatment-outcome pair, $a^{\top}Y$ and $d^{\top}T$, let $U_{a,d}$ represent a scalar projection of $U$ which captures all confounding for that treatment-outcome pair. We call $U_{a,d}$ a “minimal deconfounder” for treatment-outcome pair $a^{\top}Y$ and $d^{\top}T$ since it is the lowest dimensional function of $U$ that is sufficient to address confounding (D’Amour & Franks, 2021). There is no unique minimal deconfounder: for example, both $d^{\top}BU$ and $a^{\top}\Gamma\Sigma_{u\mid t}^{-1/2}U$ are different scalar summaries of $U$ that are sufficient to remove all confounding bias. Despite this, we can still reason about the partial variances explained by such a minimal deconfounder. Here, we consider the fraction of outcome variance explained by the minimal deconfounder given the treatment, $R^{2}_{a^{\top}Y\sim U_{a,d}|T}$, and the fraction of treatment variance explained by the minimal deconfounder, $R^{2}_{d^{\top}T\sim U_{a,d}}$. These are the sensitivity parameters one would specify in a “one-at-a-time” sensitivity analysis, e.g., using the R-squared style sensitivity analysis proposed by Cinelli & Hazlett (2020). Here, we can rewrite the bias for a given estimand in terms of the partial R-squared values on a minimal sufficient confounder as
@@
These upper bounds, when they are achieved, correspond to the worst-case bias bound in Theorem 1. One advantage of the multi-treatment, multi-outcome with factor confounding is that these upper bounds are identifiable from the observed data. Other approaches, such as the extreme robustness value approach proposed by Cinelli & Hazlett (2020), consider the implications for inference when the partial correlation between the unmeasured confounder and outcome is one (e.g., $R^{2}_{a^{\top}Y\sim U_{a,d}|T}=1$). In our setting, when we infer $\frac{a^{\top}\Gamma\Gamma^{\top}a}{a^{\top}\Sigma_{y|t}a}\ll 1$, we produce a much less conservative characterization of the range of plausible casual effects.
In practice, it may be reasonable to assume that the minimal sufficient confounder explains an even smaller fraction of the variability in the treatment and outcome than those in (11). Commonly, calibration for these R-squared variables is done by leveraging observed covariates to compute similar benchmark R-squared values. With benchmark values in hand, one could provide bounds on the bias of the causal effect, since the bias of our estimand can be written in terms of these parameters. Finally, as with bounding the treatment effects directly, joint R-squared style sensitivity analysis for multiple estimands can have large benefits relative to a one-at-a-time sensitivity analysis. The partial R-squared values for each estimand are not variationally independent, as they depend on the common parameter $R$, and thus if we require marginal bounds to be satisfied concurrently for all estimands, then the implied joint bounds for individual estimands can be much tighter. To find the partial identification regions, we can use the same sampling algorithm discussed in Section 4.1, where we retain rotation matrices that satisfy all of the partial R-squared constraints.
@@ 4.3 Negative controls
Negative controls have been widely used in observational studies to detect or mitigate bias in the causal effect of the treatment on the outcome. In the multi-treatment, multi-outcome setting, we refer to a negative control pair as any treatment and outcome pair among our $k$ treatments and $q$ outcomes for which we assume a priori that the PATE for that treatment-outcome pair is zero. Such negative control assumptions can be viewed as a special case of a sensitivity bound on the treatment effects described in Section 4.1, where the constraints in (10) are set so that $z_{l}=z_{u}=0.$ Given the importance of null controls in the literature, we elaborate on this special case below.
@@
The partial identification regions with null control constraints can be complex, even consisting of either disjoint intervals or isolated points. Below, we introduce a constrained optimization problem for deriving these partial identification regions under arbitrary negative control constraints, and derive analytic partial identification regions that are conservative, but provide intuition about the benefits of negative controls.
4.3.1 Approaches to finding negative control partial identification region
@@
and $\mathcal{V}_{m,m}$ is the Stiefel manifold of all $m\times m$ orthogonal matrices. If the bias value $\beta$ is in the partial identification region established by the negative control constraints, then $\nu_{\beta}=0$. Therefore, for any $\beta$, we minimize equation (15) over the Stiefel manifold using the optimization algorithm implemented in the rstiefel package (Hoff & Franks, 2019). To find the partial identification region, we minimize (15) over a grid of all possible values of $\beta$ and keep track of those that minimize this objective function at zero. In practice, the numerical solutions $\nu_{\beta}$ are never exactly zero, and thus we simply check that both $(a_{1}^{\top}\widetilde{\Gamma}\widetilde{R}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}-\beta)^{2}\leq\delta_{1}$ and $\sum_{j=1}^{J}\|b_{j}^{\top}\widetilde{\Gamma}\widetilde{R}M_{j}-b_{j}^{\top}\check{\mathcal{G}}_{j}\|_{2}^{2}\leq\delta_{2}$ for sufficiently small constants $\delta_{1},\delta_{2}>0$. A full description of the algorithm for identifying feasible biases can be found in Appendix B.
In addition to finding the bounds computationally, we are able to derive partial identification regions analytically. While these regions are conservative, the analytic results provide insight into the types of negative controls that lead to the largest reduction in the size of the partial identification region.
@@
Assume models (1)–(3) and Assumption 4, and suppose we only have $J=1$ negative control outcome with $c_{1}\leq m$ negative control contrasts. Let $k_{aa}=a^{\top}\Gamma\Gamma^{\top}a$, $k_{ab}=a^{\top}\Gamma\Gamma^{\top}b_{1}$, and $k_{bb}=b_{1}^{\top}\Gamma\Gamma^{\top}b_{1}$, which are all identifiable scalars from our factor model assumptions. Lastly, let $M_{1}^{\dagger}$ be a generalized inverse of $M_{1}$. Then, $\text{Bias}_{a,t_{1},t_{2}}$ is in the following interval
@@
The proof of Theorem 3 can be seen in Appendix A.3. This result shows that negative controls provide the largest benefit when (i) $a^{\top}\Gamma\Gamma^{\top}b_{1}$ is large in magnitude, (ii) the confounding bias of the negative control given by $b_{1}^{\top}\check{\mathcal{G}}_{j}$ is large, and (iii) when $M_{1}^{\dagger}$ is colinear with $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$. The first condition occurs when the negative control outcome and the outcome of interest have similar confounding mechanisms, meaning that the same linear combinations of the unmeasured confounders affect their outcomes. The second condition simply implies that the negative control estimand also suffers from bias due to the unmeasured confounders. The third condition occurs when the linear combinations of the unmeasured confounders that affect both the negative control treatment and the treatment of interest are similar. Therefore, ideal negative control pairs are those with large confounding biases and similar confounding mechanisms as the estimand of interest. While we focused on $J=1$ in Theorem 3 to build intuition, we extend this result to settings with $J>1$ in Appendix A.4.

### v3

@@ 5 Joint Partial identification regions under additional constraints
Theorem 2 highlights that with a factor confounding assumption, we can bound the causal effects for all estimands in (7) in settings with multiple treatments and multiple outcomes. However, in some contexts, these bounds are too wide to be practically useful without additional assumptions. This occurs in our motivating example in Section 7 as the bounds contain all plausible (and some implausible) effects of air pollution. However, there are additional benefits that come from considering the joint partial identification region for multiple estimands. Importantly, the joint partial identification region is not simply the Cartesian product of the marginal bounds, as the worst-case bounds for multiple estimands cannot be simultaneously attained. In many instances, assumptions about the bias on one estimand can be highly informative about the bias of other estimands (see Figure 1).
By way of example, suppose that we are interested in two estimands, $\text{PATE}_{a_{1},t_{1},t_{2}}$ and $\text{PATE}_{a_{2},t_{1},t_{2}}$, which correspond to distinct outcomes, $a_{1}$ and $a_{2}$, but the same treatment contrast ($t_{2}$ vs $t_{1}$). The degree to which assumptions about $\text{PATE}_{a_{1},t_{1},t_{2}}$ inform the partial identification region for $\text{PATE}_{a_{2},t_{1},t_{2}}$ is driven in part by the inner product of $a_{1}^{\top}\Gamma$ and $a_{2}^{\top}\Gamma$. This is illustrated in Figure 1 where we can see a strong dependence in the bivariate partial identification region when $|\langle a_{1}^{\top}\Gamma$, $a_{2}^{\top}\Gamma\rangle|$ is large (Figure 1, right). Even when $a_{1}^{\top}\Gamma$ and $a_{2}^{\top}\Gamma$ are orthogonal, the magnitude of the bias for one estimand is weakly informative about the magnitude of the bias for the other (Figure 1, left). We formalize these ideas in the following proposition, which characterizes the partial identification region for one estimand as a function of the bias of another.
Proposition 1.
Given Assumptions 14, if we further assume that $\text{Bias}_{a_{1},t_{1},t_{2}}=\beta_{1}$, then the partial identification region for a different outcome with the same treatment contrast, $\text{Bias}_{a_{2},t_{1},t_{2}}$, is given by
$\text{Bias}_{a_{2},t_{1},t_{2}}$$\in$
$\beta_{1}\left(\frac{a_{1}^{\top}\Gamma\Gamma^{\top}a_{2}}{a_{1}^{\top}\Gamma\Gamma^{\top}a_{1}}\right)$$\pm\sqrt{a_{2}^{\top}\Gamma\Gamma^{\top}a_{2}-\frac{(a_{1}^{\top}\Gamma\Gamma^{\top}a_{2})^{2}}{a_{1}^{\top}\Gamma\Gamma^{\top}a_{1}}}\left\|\left(I-\frac{\Gamma^{\top}a_{1}a_{1}^{\top}\Gamma}{a_{1}^{\top}\Gamma\Gamma^{\top}a_{1}}\right)\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\right\|_{2}$ (11)
$\subseteq\left[-\|a_{2}^{\top}\Gamma\|_{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2},~{}~{}~{}\|a_{2}^{\top}\Gamma\|_{2}\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}\right]$ (12)
This conditional partial identification region (11) is contained in the marginal identification region (12), can be much shorter in length, and may have a different center. For one, given $\text{Bias}_{a_{1},t_{1},t_{2}}=\beta_{1}$, the identification region for $\text{Bias}_{a_{2},t_{1},t_{2}}$ is no longer necessarily centered at zero. The magnitude of the bias correction depends on $\beta_{1}$ as well as $\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle$. The width of the partial identification region in (11) is the product of two terms. The first term is smallest when $|\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle|$ is large and reaches zero when $a_{1}^{\top}\Gamma$ and $a_{2}^{\top}\Gamma$ are collinear. In the air pollution setting, it is plausible that $|\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle|$ is large, as common unmeasured confounders are likely to affect multiple public health outcomes. The second term is small when $\text{Bias}_{a_{1},t_{1},t_{2}}$ is large in magnitude. This is because $\bigg{(}I-\frac{\Gamma^{\top}a_{1}a_{1}^{\top}\Gamma}{a_{1}^{\top}\Gamma\Gamma^{\top}a_{1}}\bigg{)}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$ is the projection of $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$ onto the space orthogonal to $a_{1}^{\top}\Gamma$, which is smaller when the bias, $a_{1}^{\top}\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$, is large.
While we have focused on two estimands with the same exposure contrast and different outcomes, similar intuition holds for two estimands with the same outcome and different exposure contrasts. If two estimands have different treatment contrasts, then they will have different values of the $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$ term in their respective biases, and the inner product between these vectors dictates the strength of dependence in the partial identification region. We don’t explicitly write down these partial identification regions, as they can be seen as special cases of the partial identification regions we derive later in Section 5.3 that hold in more general settings for any number of estimands.
@@
Figure 1: Comparison of bivariate partial identification regions for the bias when either $\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle=0$ or $\langle a_{1}^{\top}\Gamma,a_{2}^{\top}\Gamma\rangle<0$. The horizontal dashed lines in the left panel correspond to the marginal partial identification region for the bias of the second estimand. The horizontal dashed lines in the right panel correspond to the conditional partial identification region for the bias of the second estimand for a particular choice of the bias of the first estimand.
@@ 5.1 Sensitivity Bounds on treatment effects
In many applications, prior knowledge on a reasonable range of effect sizes is known and can be incorporated into the partial identification regions. For instance, in our application to air pollution (Section 7), we know that it is biologically implausible for pollution to have strong protective effects on health outcomes. We may also have strong prior knowledge about the largest plausible harmful effect, as it is not expected for air pollution to be the sole driver of complex public health outcomes. In our notation, a direct assumption about the magnitude of the causal effect on a single estimand takes the form
@@
If we make these assumptions for a number of estimands, we gain more information about $R$, and hence the bias for all other estimands. Analytic characterizations of the partial identification region under these constraints is a non-trivial problem and as such, we find approximate solutions using a rejection sampling algorithm to explore the range of possible causal effects satisfying (13). Specifically, we draw orthogonal matrices $R^{(h)}$ for $h=1,\dots,H$ uniformly on the Stiefel manifold of $m\times m$ orthogonal matrices, for large values of $H$. For each orthogonal matrix, $R^{(h)}$, we check whether the constraints in (13) are satisfied. Each matrix $R^{(h)}$ that satisfies the aforementioned criteria determines the bias for each estimand of interest, and we can explore the set of these matrices to determine plausible biases, and therefore partial identification regions. Since $m$ is generally small (e.g., single a digit value), such a Monte Carlo rejection sampling strategy is feasible.
@@ 5.2 Sensitivity Bounds on Partial R-squared Values
As an alternative to specifying sensitivity bounds directly on the treatment effects, we can instead specify bounds on the partial variance explained by confounders. In a single-treatment single-outcome analysis, Cinelli and Hazlett, (2020) propose specifying upper bounds on both the partial fraction of outcome variance explained by confounders ($R^{2}_{Y\sim U|T}$) and the fraction of treatment variance explained by confounders ($R^{2}_{T\sim U})$. Here, we show how their approach can be incorporated into our multi-treatment multi-outcome setting. Throughout this section, we let $d=(t_{1}-t_{2})$ correspond to the treatment contrast of interest.
For a given treatment-outcome pair, $a^{\top}Y$ and $d^{\top}T$, let $U_{a,d}$ represent a scalar projection of $U$ which captures all confounding for that treatment-outcome pair. We call $U_{a,d}$ a “minimal deconfounder” for treatment-outcome pair $a^{\top}Y$ and $d^{\top}T$ since it is the lowest dimensional function of $U$ that is sufficient to address confounding (D’Amour and Franks,, 2021). There is no unique minimal deconfounder: for example, both $d^{\top}BU$ and $a^{\top}\Gamma\Sigma_{u\mid t}^{-1/2}U$ are different scalar summaries of $U$ that are sufficient to remove all confounding bias. Despite this, we can still reason about the partial variances explained by such a minimal deconfounder. Here, we consider the fraction of outcome variance explained by the minimal deconfounder given the treatment, $R^{2}_{a^{\top}Y\sim U_{a,d}|T}$, and the fraction of treatment variance explained by the minimal deconfounder, $R^{2}_{d^{\top}T\sim U_{a,d}}$. These are the sensitivity parameters one would specify in a “one-at-a-time” sensitivity analysis, e.g., using the R-squared style sensitivity analysis proposed by Cinelli and Hazlett, (2020). Here, we can rewrite the bias for a given estimand in terms of the partial R-squared values on a minimal sufficient confounder as
@@
These upper bounds, when they are achieved, correspond to the worst-case bias bound in Theorem 1. One advantage of the multi-treatment, multi-outcome setting with factor confounding is that these upper bounds are identifiable from the observed data. Other approaches, such as the extreme robustness value approach proposed by Cinelli and Hazlett, (2020), consider the implications for inference when the partial correlation between the unmeasured confounder and outcome is one (e.g., $R^{2}_{a^{\top}Y\sim U_{a,d}|T}=1$). In our setting, when we infer $\frac{a^{\top}\Gamma\Gamma^{\top}a}{a^{\top}\Sigma_{y|t}a}\ll 1$, we produce a much less conservative characterization of the range of plausible casual effects.
In practice, it may be reasonable to assume that the minimal sufficient confounder explains an even smaller fraction of the variability in the treatment and outcome than those in (14). Commonly, calibration for these R-squared variables is done by leveraging observed covariates to compute similar benchmark R-squared values. With benchmark values in hand, one could provide bounds on the bias of the causal effect, since the bias of our estimand can be written in terms of these parameters. Finally, as with bounding the treatment effects directly, joint R-squared style sensitivity analysis for multiple estimands can have large benefits relative to a one-at-a-time sensitivity analysis. The partial R-squared values for each estimand are not variationally independent, as they depend on the common rotation matrix parameter $R$, and thus if we require marginal bounds to be satisfied concurrently for all estimands, then the implied joint bounds for individual estimands can be much tighter. To find the partial identification regions, we can use the same sampling algorithm discussed in Section 5.1, where we retain rotation matrices that satisfy all of the partial R-squared constraints.
@@ 5.3 Negative controls
Negative controls have been widely used in observational studies to detect or mitigate bias in the causal effect of the treatment on the outcome. In air pollution epidemiology, examples include future air pollutant levels or air pollutant levels in a distant area being used as negative control exposures (Lumley and Sheppard,, 2000), and appendicitis hospitalizations being used as a negative control outcome (Sheppard et al.,, 1999). Generally, in the multi-treatment, multi-outcome setting, we refer to a negative control pair as any treatment and outcome pair among our $k$ treatments and $q$ outcomes for which we assume a priori that the PATE for that treatment-outcome pair is zero. Such negative control assumptions can be viewed as a special case of a sensitivity bound on the treatment effects described in Section 5.1, where the constraints in (13) are set so that $z_{l}=z_{u}=0.$ Given the importance of null controls in the literature, we elaborate on this special case below.
@@
The partial identification regions with null control constraints can be complex, even consisting of either disjoint intervals or isolated points. Below, we introduce a constrained optimization problem for deriving these partial identification regions under arbitrary negative control constraints, and derive analytic partial identification regions that are conservative at times, but provide intuition about the benefits of negative controls.
5.3.1 Approaches to finding negative control partial identification regions
@@
and $\mathcal{V}_{m,m}$ is the Stiefel manifold of all $m\times m$ orthogonal matrices. If the bias value $\beta$ is in the partial identification region established by the negative control constraints, then $\nu_{\beta}=0$. Therefore, for any $\beta$, we minimize equation (18) over the Stiefel manifold using the optimization algorithm implemented in the rstiefel package (Hoff and Franks,, 2019). To find the partial identification region, we minimize (18) over a grid of all possible values of $\beta$ and keep track of those that minimize this objective function at zero. In practice, the numerical solutions $\nu_{\beta}$ are never exactly zero, and thus we simply check that both $(a_{1}^{\top}\widetilde{\Gamma}\widetilde{R}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}-\beta)^{2}\leq\delta_{1}$ and $\sum_{j=1}^{J}\|b_{j}^{\top}\widetilde{\Gamma}\widetilde{R}M_{j}-b_{j}^{\top}\check{\mathcal{G}}_{j}\|_{2}^{2}\leq\delta_{2}$ for sufficiently small constants $\delta_{1},\delta_{2}>0$. A full description of the algorithm for identifying feasible biases can be found in Appendix B.
In addition to finding the bounds computationally, we are able to derive partial identification regions analytically. While these regions can be conservative in certain situations, they provide useful insight into the types of negative controls that lead to the largest reduction in the size of the partial identification region.
@@
Suppose we only have $J=1$ negative control outcome with $c_{1}\leq m$ negative control contrasts and that Assumptions 14 hold. Let $k_{aa}=a^{\top}\Gamma\Gamma^{\top}a$, $k_{ab}=a^{\top}\Gamma\Gamma^{\top}b_{1}$, and $k_{bb}=b_{1}^{\top}\Gamma\Gamma^{\top}b_{1}$, which are all identifiable scalars from our factor model assumptions. Lastly, let $M_{1}^{\dagger}$ be a generalized inverse of $M_{1}$. Then, $\text{Bias}_{a,t_{1},t_{2}}$ is in the following interval
@@
The proof of Theorem 3 can be seen in Appendix A. This result shows that negative controls provide the largest benefit when (i) $a^{\top}\Gamma\Gamma^{\top}b_{1}$ is large in magnitude, (ii) the confounding bias of the negative control given by $b_{1}^{\top}\check{\mathcal{G}}_{j}$ is large, and (iii) when $M_{1}^{\dagger}$ is colinear with $\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$. The first condition occurs when the negative control outcome and the outcome of interest have similar confounding mechanisms, meaning that the same linear combinations of the unmeasured confounders affect their outcomes. The second condition simply implies that the negative control estimand also suffers from bias due to the unmeasured confounders. The third condition occurs when the linear combinations of the unmeasured confounders that affect both the negative control treatment and the treatment of interest are similar. Therefore, ideal negative control pairs are those with large confounding biases and similar confounding mechanisms as the estimand of interest. While we focused on $J=1$ in Theorem 3 to build intuition, we extend this result to settings with $J>1$ in Appendix A.